OpenStudy (abbycross167):

Can someone please check my math questions please? 42. Solve r/5 – 6 = -1. a. r = 1 b. r = 5 c. r = 10 d. r = 25 My Answer~ d 43. The height of an ostrich is 20 inches more than 4 times the height of a kiwi. If an ostrich is 108 inches tall, how tall is a kiwi? a. 22 inches b. 27 inches c. 44 inches d. 66 inches My Answer~b

OpenStudy (anonymous):

42. Correct 43. Incorrect Try again and if you get stuck let me know :)

OpenStudy (abbycross167):

ok, do you think 43 is a? I did the math over and got 'a' as the answer

OpenStudy (abbycross167):

@CShrix

OpenStudy (phi):

y=20+4x (y is ostrich, x is kiwi) 20+4x= 108 4x= 108-20 4x= 88 x= 88/4 x=22

OpenStudy (abbycross167):

Alright! Thank you very much!!